OpenStudy (anonymous):

Please hellpppp medal rewarded If a bird can fly 30 miles with the wind in the same time it can fly 18 miles against the wind when the wind speed is 8 mph whats the speed of the bird in still air

OpenStudy (anonymous):

Wait this question is weird...

OpenStudy (anonymous):

Yea the book says its 3

OpenStudy (anonymous):

Okay hold on...

OpenStudy (anonymous):

If a bird can fly 30 miles with the wind in the same time it can fly 18 miles against the wind when the wind speed is 8 mph \[ 30/(b+8) = t = 18/(b-8) \]This comes from: \[ v = d/t\implies t = d/v \]

OpenStudy (anonymous):

It simplifies to \[ 30(b-8) = 18(b+8) \]

OpenStudy (anonymous):

@natth88 Now try.

OpenStudy (anonymous):

I did that before and got thesamething32

OpenStudy (anonymous):

It's ok ill just ask my teacher tomorrow thanks

OpenStudy (anonymous):

You got 32?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Okay that is really bizarre. When you figure it out tell me.